OpenStudy (anonymous):

can you factor this out for me: 2x^6 + 2sqrt(6)x^3 + 3

OpenStudy (anonymous):

2u^2 + 2sqrt(6) u + 3

OpenStudy (anonymous):

\[u= \frac{ - 2\sqrt{6} +- \sqrt{ 24 -4(2)(3) } } { 4} \]

OpenStudy (anonymous):

wghere u=x^3

OpenStudy (anonymous):

\[u = \frac{ -\sqrt{6} } { 2} \]

OpenStudy (anonymous):

duno understand

OpenStudy (anonymous):

\[= (x^3 + \frac{ \sqrt{6} } { 2} ) ^2 \]

OpenStudy (anonymous):

Im supposed to factor this without that formula. By direct factorization. Please help

OpenStudy (anonymous):

nah thiik something wrong

OpenStudy (anonymous):

Im trhing to help a friend and I cannot find the answer. Thanks